# Polarimeter to Unify the Corona and Heliosphere (PUNCH)

### v1

@@ 2.2.2 Science Topic 2B: Stream Interaction Regions
Prior to STEREO/HI, SIR research mainly depended on radio interplanetary scintillation (IPS) observations (Breen et al., 1998; Bisi et al., 2010; Dorrian et al., 2010) or used in-situ measurements that typically sample a single track through the feature (Tappin, Hewish, and Gapper, 1984; Pizzo and Gosling, 1994; Möbius et al., 2002; Bucik et al., 2009). From the wide-angle view on SIRs provided by heliospheric imagers, we now know SIRs manifest very differently depending on whether one observes from ahead looking east, or from behind looking west (Wood et al., 2010). In particular, the east view observes the advancing SIR structure along its developed part, far from the Sun, while the west view observes along a much closer, earlier phase of evolution during the short interval as the SIR sweeps past (Figure 6).
@@ 2.2.3 Science Topic 2C: Shocks
The basic geometric theory of interplanetary shocks is well established (Hundhausen, 1985), but large-scale shock structure is rarely simple. Shock morphology depends on the surrounding medium (Odstrcil and Pizzo, 2009; Tappin and Simnett, 1997). Many shocks and CME sheaths appear to have an asymmetric structure (Pizzo et al., 2011; Feng et al., 2012). Further, some shocks appear to form on the CME flanks rather than, as expected, at its nose where density pileup is greatest (Ontiveros and Vourlidas, 2009; Wood et al., 2009; Wood, Howard, and Socker, 2010).
@@
Prior generations of coronagraphs and heliospheric imagers have been used to explore shocks, as well as the interaction of CMEs with each other and with other large-scale structures such as SIRs (Lugaz et al., 2012; Liu et al., 2012; Temmer et al., 2012; Harrison et al., 2012; Möstl et al., 2012; Wood et al., 2012). 3D analyses using the Solar Mass Ejection Imager (SMEI) data (Jackson et al., 2009; Jackson et al., 2011) show that shocks appear “spotty,” reminiscent of expectations for turbulent instabilities. However, such studies have been severely hampered by limitations associated with instrument sensitivity and motion blur.
@@
3 Observing Requirements
Each topic in Section 2 levies specific observing requirements on the PUNCH mission. During mission definition, we developed a set of mission-wide requirements that encompasses all these specific observing requirements. Here we describe the seven primary mission-wide observing requirements (“Baseline Science Mission Requirements,” or BSRs) and the questions that drive them. These requirements flow down to the mission definition, and formed the basis of the engineering requirements structure for all aspects of PUNCH.
In many cases, an observing requirement varies across the field of view. For example, spatial resolution can be coarser far from the Sun than close to it, because of the known overall expansion of streams and structures in the solar wind. In these cases, values are specified at two representative radial positions: 15 R (3.75°) from Sun center, to capture the outer corona in the vicinity of the Alfvén zone and within the region visually dominated by radial striae, and 80 R (20°) from Sun center, to capture the young solar wind above the transition from anisotropic to isotropic structure (DeForest et al., 2016).
@@ 3.1 Field of View and Duration
The mission duration is set by the need to observe ten well-presented large CMEs over the course of the mission lifetime (Section 2.2.1). Based on a launch into the active phases of the solar cycle, existing counts in the CaCTUS catalog (Robbrecht, Berghmans, and Van der Linden, 2009), and a selection factor of 2.5% to account for CME size, presentation, direction of propagation, overlap, and lunar interference, we estimated 0.025 well-presented, well separated, front-side “PUNCH large CME events” per day, driving a mission duration requirement of 15 months and (with margin) the overall mission design duration of 24 months.
@@ 3.3 Polarization
The relationship between photometry of individual camera images through a polarizing filter, and polarimetry using multiple images through a trio of such filters, is nontrivial but tractable via algebraic analysis. DeForest, Seaton, and West (2022) recently revisited that relationship via derivation from basic polarization physics. As a rule of thumb: noise levels in PUNCH unpolarized $B$ data products (derived from triplets of images through polarizing filters) scale approximately as would those in sums of unpolarized exposures; while noise levels in $pB$ data products from PUNCH are worse by a factor of at most $\sqrt{3}$, compared to the corresponding $B$ data product.
@@ 4.2 Spacecraft design
The thermal control system is a cold-biased passive design with supplementary multi-zone digitally controlled management heaters and thermostatically switched survival heaters. Spacecraft subsystems dump heat into the body of the spacecraft. The instruments are thermally isolated from the spacecraft, and radiate their waste heat to space. The instruments include CCD cameras that are passively cooled by radiation to deep space; this takes advantage of the roll attitude program for WFI, and drives a similar roll program for NFI, both of which maintain the CCD radiator in shade on the zenith side of the spacecraft bus.
@@ 5 PUNCH instruments
PUNCH operates as a single “virtual instrument” built from three individual physical instruments. The Narrow Field Imager (NFI; Section 5.1) and three Wide Field Imagers (WFIs; Section 5.2) are physically matched such that each instrument’s instantaneous field of view (IFOV) overlaps with each of the other three instruments’ IFOVs; the wavelength passbands are matched at 450–750 nm; and the observing sequences are synchronized to make the merged science data as seamless as possible.
@@ 5.1 Narrow Field Imager
The Narrow Field Imager (NFI) is an externally-occulted compact coronagraph design built by the U.S. Naval Research Laboratory (NRL). NFI is diagrammed in Figure 10. The instrument is described in detail by (Colaninno et al., 2025); the design is summarized here. NFI builds on a long history of coronagraph development, beginning with Lyot (1930) and extending through OSO-7 (Koomen et al., 1975), Skylab-ATM (MacQueen et al., 1974), Solwind (Sheeley et al., 1980), SOHO/LASCO (Brueckner et al., 1995), STEREO/SECCHI (Howard et al., 2008), and GOES/CCOR (Thernisien et al., 2021; Dudley et al., 2023).
@@
A front-end baffle assembly (green) supports an A0 aperture with a toothed outer edge that minimizes edge diffraction into the optics, and a pylon and occulter assembly that block direct sunlight from entering the optics. The occulter is a tapered-ogive [^5] Readers are reminded that an “ogive” is a figure of revolution of a circle, about a non-diameter chord of that circle; or a truncated segment of such a figure. multi-disk design with a rear-facing conical light trap. The multiple disks force multiple diffraction scatter events, strongly attenuating diffracted sunlight. The conical light trap absorbs scattered light from the rear of the baffle assembly. Most sunlight entering the instrument impacts a spherical Heat Rejection Mirror (HRM) at the back of the baffle assembly, and is focused into the empty space opposite the pylon, and ejected into space. The HRM has a central hole that forms the aperture to a lens barrel assembly that focuses light through a polarizing filter wheel (Section 5.4) onto a CCD detector connected to an interface camera module, and operated by an external Camera Electronics Box (Section 5.3). The entire instrument was closed out, at launch, by a ground-resettable one-time-open door. The CCD is cooled by thermal conduction through a thermal strap to a radiator aimed at deep space. The entire instrument structure is supported on the spacecraft by a kinematic mount comprising three feet that bolt directly to the spacecraft.
@@
During early operations, stray light in the NFI instrument was found to be greater, and far more variable, than anticipated during design and development. This has been attributed to variable Earthshine illuminating the interior of the baffle tube, and thereby scattering onto the rear of the occulter. The process to scrub this variable stray light from the data is the subject of a rapid development effort early in the mission; it is described in more detail by Hughes et al. (2025).
@@ 5.2 Wide Field Imager
The Wide Field Imager (WFI) is a corral-baffle heliospheric imaging design built at Southwest Research Institute (SwRI). WFI is diagrammed in Figure 11. The instrument is described in detail by (Laurent et al., 2025); the design is summarized here. WFI is analogous to an externally-occulted coronagraph, in linear geometry (i.e. the baffle structure is approximately symmetric via extrusion along the spacecraft Y axis, rather than via revolution about the spacecraft X axis). Heliospheric imaging was conceived in the late 20th century using the three photometers of the spinning Helios mission (Jackson and Leinert, 1985), and developed through Coriolis/SMEI (Eyles et al., 2003; Jackson et al., 2004), STEREO/HI (Howard et al., 2008; Eyles et al., 2009), SolO/SOLOHI (Howard et al., 2020), and PSP/WISPR (Vourlidas et al., 2016). The more modern designs include a full optical imaging system, a solar baffle, and a deeply vaned light trap to absorb glint; WFI is similar, with the addition of a lunar baffle near the optical aperture, to reduce flooding by moonlight.
@@ 5.3 Camera & onboard data handling
Camera pixels are square-root coded for compression: the digitizer value $P_{xy}$ of each pixel is multiplied by an integer scale factor $\alpha$ to a total of up to 32 bits width, then the integer square root of that value, $(\alpha P_{xy})^{0.5}$, is recorded instead of $P_{xy}$. The square root operation reduces the pixel values back to 16 bit integers, which typically occupy fewer bits of actual dynamic range. The purpose of the square root operation is to match digital transition step size in the final data product, to the variable level of quantum shot noise across the dynamic range (DeForest et al., 2022). This in turn enables further compression of the remaining bits of each image, by removing useless high-entropy noise while preserving useful and lower-entropy data. In flight, $\alpha$ was selected to be 16, or $2^{4}$. WFI images, single 16-bit-deep camera frames, thus retain only 10 bits of information (roughly 1,000 unique values) across the range of the digitizer. NFI raw images, being sums of three camera frames, begin as 18-bit integer values occupying 17.5 bits of dynamic range, and conclude with 10.75 bits of dynamic range (roughly 1700 unique values). This setting preserves a small amount of photon noise to dither photometric values, so that digital steps do not dominate the images. This dynamic-range “toeroom” (room to expand at the bottom) is at least 0.7 (RMS) digitizer counts at all brightness levels, and 1.5 (RMS) digitizer counts across most of the dynamic range. The smaller values occur for dark pixels, where the camera digitizer “pedestal” (approximately 400 counts per exposure) is significant compared to the photon signal; and the larger values occur at pixel values above roughly 2,000 counts (or 6,000 counts for NFI sums) where photon counting noise is the most important element of uncertainty in each pixel value.

### v2

@@ 2.2.2 Science Topic 2B: Stream Interaction Regions
Prior to STEREO/HI, SIR research mainly depended on radio interplanetary scintillation (IPS) observations (Breen et al., 1998; Bisi et al., 2010; Dorrian et al., 2010) or used in-situ measurements that typically sample a single track through the feature (Tappin et al., 1984; Pizzo and Gosling, 1994; Möbius et al., 2002; Bucik et al., 2009). From the wide-angle view on SIRs provided by heliospheric imagers, we now know SIRs manifest very differently depending on whether one observes from ahead looking east, or from behind looking west (Wood et al., 2010b). In particular, the east view observes the advancing SIR structure along its developed part, far from the Sun, while the west view observes along a much closer, earlier phase of evolution during the short interval as the SIR sweeps past (Figure 6).
@@ 2.2.3 Science Topic 2C: Shocks
The basic geometric theory of interplanetary shocks is well established (Hundhausen, 1985), but large-scale shock structure is rarely simple. Shock morphology depends on the surrounding medium (Odstrcil and Pizzo, 2009; Tappin and Simnett, 1997). Many shocks and CME sheaths appear to have an asymmetric structure (Pizzo et al., 2011; Feng et al., 2012). Further, some shocks appear to form on the CME flanks rather than, as expected, at its nose where density pileup is greatest (Ontiveros and Vourlidas, 2009; Wood et al., 2009, 2010a).
@@
Prior generations of coronagraphs and heliospheric imagers have been used to explore shocks, as well as the interaction of CMEs with each other and with other large-scale structures such as SIRs (Lugaz et al., 2012; Liu et al., 2012; Temmer et al., 2012; Harrison et al., 2012; Möstl et al., 2012; Wood et al., 2012). 3D analyses using the Solar Mass Ejection Imager (SMEI) data (Jackson et al., 2009, 2011) show that shocks appear “spotty,” reminiscent of expectations for turbulent instabilities. However, such studies have been severely hampered by limitations associated with instrument sensitivity and motion blur.
@@
3 Science Requirements
Engineering requirements are central to development of any mission, because large projects in general and space missions in particular represent trades between capability and cost. Because the process to develop mission-level requirements is not as familiar in the scientific literature as reporting of the resulting instrument performance, we describe here both the PUNCH top-level observing requirements and an overview of the methodology used to develop them.
Each topic in Section 2 levies specific scientific observing requirements on the PUNCH mission: aspects of observation that are needed to resolve the corresponding science question. These requirements, such as spatial/angular resolution and photometric sensitivity, are specific to each topic and were derived from review of the scientific literature at the start of mission definition. Distilling these scientific requirements into testable engineering requirements meant establishing a complete minimum set of mission parameters to constrain, rendering all the individual constraints on those parameters into a form that is directly comparable across questions, and identifying a notional set of “Baseline Science Requirements” (BSRs) that would satisfy all of the science questions. These requirements became the mission definition. We executed the full derivation process twice, using two different criteria. We developed the BSRs using best-estimate observing requirements to fully address each particular scientific question, with an identified analysis path to closure of the question. We also developed a separate set of “Threshold Science Requirements” (TSRs) based on best-estimate observing requirements to significantly advance each scientific question. The BSRs were used by the instrument and engineering teams to derive a series of lower level requirements which, in turn, drove each aspect of the mission design. The TSRs provided understanding of how flexible each requirement might be, enabling development trades or compromises in case they were needed.
Developing mission top-level requirements itself requires compromise between three motivations: first, the desire to fully constrain the capability of a new instrument suite, which determines the minimum set of requirements; second, the need for each requirement to fully verifiable during development, which forces each requirement to refer to testable properties of the mission (without reference to the science target); third, the need for requirements to be feasible, which forces simple structure and requirements that can be analyzed and tested straightforwardly.
PUNCH science is broad enough that several requirements were challenging to specify. The needed image resolution varies across the field of view and is therefore specified at two representative solar elongations. Observing cadence varies for different targets and drives overall mission design, so the requirement is more complex to avoid overspecification. Photometric sensitivity is coupled to image resolution, and therefore required a novel formulation. These nuances are more fully described in the subsections below.
@@ 3.1 Field of View and Duration
The mission duration is set by the need to observe ten well-presented large CMEs over the course of the mission lifetime (Section 2.2.1). Based on a launch into the active phases of the solar cycle, existing counts in the CaCTUS catalog (Robbrecht et al., 2009), and a selection factor of 2.5% to account for CME size, presentation, direction of propagation, overlap, and lunar interference, we estimated 0.025 well-presented, well separated, front-side “PUNCH large CME events” per day, driving a mission duration requirement of 15 months and (with margin) the overall mission design duration of 24 months.
@@ 3.3 Polarization
The relationship between photometry of individual camera images through a polarizing filter, and polarimetry using multiple images through a trio of such filters, is nontrivial but tractable via algebraic analysis. DeForest et al. (2022b) recently revisited that relationship via derivation from basic polarization physics. As a rule of thumb: noise levels in PUNCH unpolarized $B$ data products (derived from triplets of images through polarizing filters) scale approximately as would those in sums of unpolarized exposures; while noise levels in $pB$ data products from PUNCH are worse by a factor of at most $\sqrt{3}$, compared to the corresponding $B$ data product.
@@ 4.2 Spacecraft design
Figure 9: The PUNCH ADCS (Blue Canyon XACT) system exceeds performance requiremens for slew recovery, filter wheel rotation recovery, and stability during camera exposures. Left: $30^{\circ}$ roll slew and recovery. Right: performance during a typical 8-minute observing window. Polarizing filter wheel rotations, and the seven standard observing-sequence exposures (Fig. 8), are visible starting 60s into the 8-minute window.
@@
The thermal control system is a cold-biased passive design with supplementary multi-zone digitally controlled management heaters and thermostatically switched survival heaters. Spacecraft subsystems dump heat into the body of the spacecraft. The instruments are thermally isolated from the spacecraft, and radiate their waste heat to space. The instruments include CCD cameras that are passively cooled by radiation to deep space; this takes advantage of the roll attitude program for WFI, and drives a similar roll program for NFI, both of which maintain the CCD radiator in shade on the +Z (zenith) side of the spacecraft bus.
@@ 5 PUNCH instruments
PUNCH operates as a single “virtual instrument” built from four individual physical instruments. The Narrow Field Imager (NFI; Section 5.1) and three Wide Field Imagers (WFIs; Section 5.2) are physically matched such that each instrument’s instantaneous field of view (IFOV) overlaps with each of the other three instruments’ IFOVs; the wavelength passbands are matched at 450–750 nm; and the observing sequences are synchronized to make the merged science data as seamless as possible.
@@ 5.1 Narrow Field Imager
The Narrow Field Imager (NFI) is an externally-occulted compact coronagraph design built by the U.S. Naval Research Laboratory (NRL). Figure 11. The instrument is a cutaway diagram of the instrument. Figure 12 is a photograph of the instrument undergoing stray light testing inside the SCOTCH vacuum chamber at NRL. NFI is described in detail by (Colaninno et al., 2025); the design is summarized here. NFI builds on a long history of coronagraph development, beginning with Lyot (1930) and extending through OSO-7 (Koomen et al., 1975), Skylab-ATM (MacQueen et al., 1974), Solwind (Sheeley et al., 1980), SOHO/LASCO (Brueckner et al., 1995), STEREO/SECCHI (Howard et al., 2008), and GOES/CCOR (Thernisien et al., 2021; Dudley et al., 2023).
@@
Figure 12: Photograph of NFI during stray light testing in NRL’s SCOTCH vacuum-testing facility shows the major front end elements diagrammed in Figure 11. The toothed aperture, single external occulter, and support pylon are brightly illuminated by a simulated solar beam; the instrument stray light housing baffle, covered in multi-layer insulation, is visible behind them. The door is open in the flight configuration at top.
In Figure 11, a front-end baffle assembly (green) supports an A0 aperture with a toothed outer edge that minimizes edge diffraction into the optics, and a pylon and occulter assembly that block direct sunlight from entering the optics. The occulter is a tapered-ogive [^5] Readers are reminded that an “ogive” is a figure of revolution of a circle, about a non-diameter chord of that circle; or a truncated segment of such a figure. multi-disk design with a rear-facing conical light trap. The multiple disks force multiple diffraction scatter events, strongly attenuating diffracted sunlight. The conical light trap absorbs scattered light from the rear of the baffle assembly. Most sunlight entering the instrument impacts a spherical Heat Rejection Mirror (HRM) at the back of the baffle assembly, and is focused into the empty space opposite the pylon, and ejected into space. The HRM has a central hole that forms the aperture to a lens barrel assembly that focuses light through a polarizing filter wheel (Section 5.4) onto a CCD detector connected to an interface camera module, and operated by an external Camera Electronics Box (Section 5.3). The entire instrument was closed out, at launch, by a ground-resettable one-time-open door. The CCD is cooled by thermal conduction through a thermal strap to a radiator aimed at deep space. The entire instrument structure is supported on the spacecraft by a kinematic mount comprising three feet that bolt directly to the spacecraft.
@@
During development NFI was tested for solar stray light rejection, at NRL’s Solar Coronagraph Optical Test Chamber (SCOTCH) facility (Korendyke, 1993). SCOTCH is a long-tube baffled vacuum chamber with a solar simulator light source, designed to measure stray light while eliminating Rayleigh and Mie scattering from the air and dust present in a more typical laboratory environment. NFI was fully illuminated by the simulated solar beam, and light levels were measured in image data collected through the instrument itself. Solar stray light in NFI is dominated by backscatter off the HRM, which illuminates the rear of the occulter. Doubly scattered sunlight enters the optics, forming an out-of-focus image of the occulter itself, with effective radiance comparable to that of the solar F corona. This anticipated effect is removed in post processing in the science reduction pipeline, described in Section 6. Solar stray light from the simulated solar beam was measured at three radial locations; these are tabulated in Table 3. The stray light was measured with each of three polarizers in the beam, and with a clear (non-polarizing) filter. The largest of the three values is reported in Table 3. The units are given as $B^{\prime}_{\odot}$ rather than $B_{\odot}$ to acknowledge linear scaling for the measured intensity of the simulated solar beam. NFI solar stray light meets design requirements at all locations.
Radial location Requirement Measured stray radiance
$6.6R_{\odot}$ $41\times 10^{-11}B^{\prime}_{\odot}$ $3.8\times 10^{-11}B^{\prime}_{\odot}$
$15R_{\odot}$ $13\times 10^{-11}B^{\prime}_{\odot}$ $5.8\times 10^{-11}B^{\prime}_{\odot}$
$27R_{\odot}$ $3.5\times 10^{-11}B^{\prime}_{\odot}$ $2.0\times 10^{-11}B^{\prime}_{\odot}$
Table 3: Solar stray light requirements and values measured for NFI at three field-of-view locations, at NRL’s SCOTCH facility
Figure 13: Early example images from NFI demonstrate instrument capability. Top: the new Moon first passed through the NFI FOV on 2025 April 27, providing a scale and brightness reference for the field of view. The Earth-illuminated Moon is approximately 2x brighter than the fixed +variable stray light pattern. Bottom: a northbound CME launched on 2025 June 2; preliminary data reduction shows that NFI resolved fine interior details of the CME as it passed through the outer corona.
During early operations, total stray light in the NFI instrument was found to be greater, and far more variable, than the solar stray light performance in Table 3. This has been attributed to variable Earthshine illuminating the interior of the baffle tube, and thereby scattering onto the rear of the occulter; the effect was not characterized during the SCOTCH testing campaign in the pre-launch phase. The additional variable stray light is up to $5\times$ brighter than the solar stray light, though more typically $0.2\times$ to $2\times$ as bright; and is confined to the zone (inside of approximately $22R_{\odot}$) covered by the defocused image of the occulter. The process to scrub this variable stray light from the data is the subject of a rapid development effort early in the mission; it is described in more detail by Hughes et al. (2025).
NFI sample data, collected during commissioning, included images of the new moon and of several bright CMEs. Example images are in Figure 13, and reveal the instrument’s high sensitivity and resolution even with preliminary data processing.
Figure 15: Photograph of WFI-2 during final instrument integration shows the major elements diagrammed in Figure 14. From front to back, the major features are the solar shade, coated in white thermal paint; the multi-vaned solar baffle, with its Acktar optical black coating; the deep light trap vanes, with Aeroglaze Z307 black coating; the square-geometry lunar baffle; and the optical aperture. The CCD radiator is protected by a protective shield. At left is the door release mechanism. The door is in its deployed flight configuration.
@@ 5.2 Wide Field Imager
The Wide Field Imager (WFI) is a corral-baffle heliospheric imager design; three were built for PUNCH at Southwest Research Institute (SwRI). Figure 14 is a diagram of WFI. Figure 15 is a photograph of one of the instruments, under development. WFI is described in detail by (Laurent et al., 2025); the design is summarized here. The is analogous to an externally-occulted coronagraph, in linear geometry (i.e. the baffle structure is approximately symmetric via extrusion along the spacecraft Y axis, rather than via revolution about the spacecraft X axis). Heliospheric imaging was conceived in the late 20th century using the three photometers of the spinning Helios mission (Jackson and Leinert, 1985), and developed through Coriolis/SMEI (Eyles et al., 2003; Jackson et al., 2004), STEREO/HI (Howard et al., 2008; Eyles et al., 2009), SolO/SOLOHI (Howard et al., 2020), and PSP/WISPR (Vourlidas et al., 2016a). The more modern designs include a full optical imaging system, a solar baffle, and a deeply vaned light trap to absorb glint; WFI is similar, with the addition of a lunar baffle near the optical aperture, to reduce flooding by moonlight.
@@
One WFI was tested for solar stray light rejection, at NRL’s SCOTCH facility. A simulated solar beam was directed at the WFI baffle, as in the NFI test, and the instrument itself was used to measure the radiance of stray light; details are given by Laurent et al. (2025). The WFI stray light rejection performance exceeded considerably the stray light rejection performance of the SCOTCH facility, and WFI images of the SCOTCH interior were dominated by interior features rather than by instrumental stray light. Typical dark chamber-interior image points measured a local radiance near $1\times 10^{-11}B^{\prime}_{\odot}$ (compared to $B^{\prime}_{\odot}$, the average radiance of the simulated solar beam in the SCOTCH chamber), and differential hit-miss beam measurements afforded sensitivity as low as a few $\times 10^{-14}B^{\prime}_{\odot}$; this was not sufficiently sensitive to measure the required FOV-average solar stray light radiance attenuation of $10^{-16}B^{\prime}_{\odot}$ directly. Instead, we used the measured lens stray-light attenuation of $3\times 10^{-3}$, a geometric factor for the apparent size of the solar vane edge “bright line” in the field of view, and differential measurement of the bright line. This yielded calculated overall solar stray light levels of a few $\times 10^{-18}B^{\prime}_{\odot}$. Given the wide margin on stray light rejection, subsequent WFIs were not end-to-end tested for stray light performance. In-flight stray light performance was found to meet requirements with wide margin, as expected.
WFI is subject to variable stray light from Earth, near both solstices and especially during the “eclipse season” surrounding the unfavorable (winter) solstice. During the eclipse season, the Earth is in front of WFI, approaching the actual instrument field of view, during passage over the north pole; and stray light increases by up to three orders of magnitude near the baffle in the corners of the image. This contaminates some images each orbit. These periods were considered during mission design and are omitted from the mission total duration count. Details are given in Laurent et al. (2025).
Figure 16 shows first light from the WFI-2 instrument, demonstrating low stray light performance and field of view. The limiting magnitude is roughly 11 in a single image: asteroid Iris, with magnitude 10.2, is clearly visible in the original data. The brightest object visible to WFI without saturation is Saiph ($\kappa$ Orionis, magnitude 2.1), which registers a peak brightness of up to 65,000 digitizer counts (of 65,535 possible) when optimally aligned with the WFI-2 pixel grid.
Figure 16: First-light image from WFI shows the very large field of view of each WFI camera, spanning over $40^{\circ}$ of sky. The dominant source of brightness is the zodiacal light (extended F corona). Several relevant constellations and objects have been marked. Asteroid Iris (magnitude 10.2) is clearly visible in the data, indicating good photometric sensitivity.
@@ 5.3 Camera & onboard data handling
Camera pixels are square-root coded for compression: the digitizer value $P_{xy}$ of each pixel is multiplied by an integer scale factor $\alpha$ to a total of up to 32 bits width, then the integer square root of that value, $(\alpha P_{xy})^{0.5}$, is recorded instead of $P_{xy}$. The square root operation reduces the pixel values back to 16 bit integers, which typically occupy fewer bits of actual dynamic range. The purpose of the square root operation is to match digital transition step size in the final data product, to the variable level of quantum shot noise across the dynamic range (DeForest et al., 2022a). This in turn enables further compression of the remaining bits of each image, by removing useless high-entropy noise while preserving useful and lower-entropy data. In flight, $\alpha$ was selected to be 16, or $2^{4}$. WFI images, single 16-bit-deep camera frames, thus retain only 10 bits of information (roughly 1,000 unique values) across the range of the digitizer. NFI raw images, being sums of three camera frames, begin as 18-bit integer values occupying 17.5 bits of dynamic range, and conclude with 10.75 bits of dynamic range (roughly 1700 unique values). This setting preserves a small amount of photon noise to dither photometric values, so that digital steps do not dominate the images. This dynamic-range “toeroom” (room to expand at the bottom) is at least 0.7 (RMS) digitizer counts at all brightness levels, and 1.5 (RMS) digitizer counts across most of the dynamic range. The smaller values occur for dark pixels, where the camera digitizer “pedestal” (approximately 400 counts per exposure) is significant compared to the photon signal; and the larger values occur at pixel values above roughly 2,000 counts (or 6,000 counts for NFI sums) where photon counting noise is the most important element of uncertainty in each pixel value.